The United States Bureau of Labor Statistics (BLS) conducts the Quarterly Census of Employment and Wages (QCEW) and reports a variety of information on each county in America. In the third quarter of 2016, the QCEW reported the total taxable earnings, in millions, of all wage earners in all 3222 counties in America.

Suppose that James is an economist who collects a simple random sample of the total taxable earnings of workers in 52 American counties during the third quarter of 2016. According to the QCEW, the true population mean and standard deviation of taxable earnings, in millions of dollars, by county are

0 0
Answer #1

Lt x be thu fotal taxable eAenn ven l28.29 6: 33 493 ns52 ) The probability tmat the mean taxable wages in James sqnmple 52PCR2=- P[2< 26-25&29 33.493/52 < ~2.29 4.645 /- = 1-P[<-0.49 = 1- 03)21 by using 2 tab le 6L89.0 - PCR>2] = O687 9